# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ Abstract
Continuous-time Conflict Based-Search (CCBS) has long been viewed as the de-facto optimal solver for multi-agent path finding in continuous time (MAPFR). Recent findings, however, show that the original theoretical variant of CCBS can suffer from non-termination, while the widely used implementation can return sub-optimal solutions. We introduce an analytical framework that yields simple and sufficient conditions under which any CCBS-style algorithm is both sound, i.e., returns only optimal solutions, and solution complete, i.e., terminates on every solvable MAPFR instance. Investigating the publicly available implementation of CCBS reveals that it violates these conditions. Though this merely indicates that CCBS might be unsound, this indication is supported by counter-examples.
@@
Leveraging the analytical framework, we propose a novel branching rule and prove that it satisfies the sufficient conditions, thereby restoring soundness and termination guarantees. Consequently, the resulting CCBS variant is both sound and solution complete, matching the guarantees of the discrete-time CBS for the first time in the continuous domain. We experimentally apply standard CCBS and CCBS under our branching rule to an example problem, with our branching rule returning a solution with lower sum-of-costs than standard CCBS. Because the branching rule largely only affects the branching step, it can be adopted as a drop-in replacement in existing code-bases, as we show in our provided implementation. Beyond CCBS, the analytical framework and termination criterion provide a systematic way to evaluate other CCBS-like MAPFR solvers and future extensions.
I INTRODUCTION
@@
As robotics and automation take an increasingly larger role in society, there is a growing need to move agents from where they are to where they need to be without collision. This is the problem of Multi-agent Path Finding (MAPF) which finds application across diverse domains, from computer games [1] to warehouses [2, 3], factories [4] to airports [5]. Progress toward fast and exact solution methods has been steady for decades, targeting the many variants of MAPF.
@@
In its classical formulation, MAPF is about finding a collision-free path for each agent on a graph from its start vertex to goal vertex. Agents are represented as points and can wait at vertices and traverse edges, with time being discrete, edges taking unit-time to traverse, and an objective function based on arrival times at goal vertices being minimized. This problem is NP-hard [6] and has many variations, comprehensively detailed in e.g. [7, 8, 9]. Early focus was placed on reducing computation and searching for sub-optimal solutions by, e.g., establishing a priority order among agents or restricting the search to a limited time-window [10, 11]. Shortly thereafter, exact methods such as $M^{*}$ [12], Increasing Cost Tree Search [13] and Conflict Based-Search (CBS) [14] were introduced. These methods cleverly avoid searching the entire solution space — which grows exponentially with the number of agents — by adopting a “lazy” approach of planning agents individually and addressing conflicts only as they occur. Of the three, CBS has enjoyed the largest adoption, with multiple continuations. To name a few, Meta-Agent CBS [15] and Improved CBS [16] for performance enhancements, Enhanced CBS [17] for bounded sub-optimal solutions, Rolling Horizon Collision Resolution [18] for lifelong MAPF, and Multi-Goal CBS [19] for MAPF where agents are assigned multiple goal vertices.
@@
The simplifying assumptions in MAPF that time is discretized and edges take unit-time to traverse bring significant drawbacks to its practical applicability; in real-world settings, environments do not always follow grid-like structures, agent speeds and action times vary, and artificial wait times to ensure that agents move in lockstep inflates the objective value. The continuous-time MAPF variant, MAPFR [20], removes such discrete-time assumptions. In MAPFR, agents can traverse edges at any real-valued time by following a motion function that defines a continuous trajectory between two connected vertices. Additionally, agents have volumes such that two agents collide when their volumes overlap. The work in [21] addresses MAPFR for sub-optimal solutions, while [22] finds optimal solutions to a simplified MAPFR problem where agents traverse edges in straight lines at constant speeds. Currently, the only known method to address MAPFR in its entirety for optimal solutions is Continuous-time CBS (CCBS) [20] which claims to be sound and solution complete, that is, it guarantees to return an optimal solution if one exists. Since the introduction of CCBS, numerous continuation works have been published, such as [23] for performance enhancements, $T$-Robust CCBS [24] for solutions that are robust to delays, [25] for Any-Angle MAPF, Continuous-time Prioritize Lifelong Planner [26] for lifelong MAPF, and [27] where CCBS is modified and applied to quadcopter drones.
@@
Despite CCBS’s wide-spread adoption, only recently has a critical assessment in [28] of the claims made in [20] established that CCBS does not fulfill its guarantees. The theoretical description of CCBS does not guarantee termination under the existence of an optimal solution (a condition for solution completeness) and the publicly available implementation does not guarantee that a returned solution is optimal (a condition for soundness). This is shown in [28] theoretically and with counter-examples. Therefore, these results imply that currently there does not exist a MAPFR solver for optimal solutions.
In this work, we develop an analytical framework to aid the analysis of CCBS and find sufficient but not necessary requirements for CCBS to be sound and solution complete. We then show that the branching rule in CCBS does not satisfy this requirement, which by itself does not mean that CCBS is not sound and solution complete, but it does complement the results in [28] and our experiments. Using the analytical framework, we then present a novel branching rule and prove that CCBS is sound and solution complete when using it. With this work, we therefore provide the first provably sound and solution-complete method to the MAPFR problem.
@@
The outline of this article is the following: In Section II, we formally define the MAPFR problem, soundness, and solution completeness. CCBS is also introduced, along with details surrounding the findings in [28]. The framework for analyzing branching rules and CCBS’s algorithmic properties is presented in Section III. In Section IV, we apply the framework on CCBS’s implemented branching rule. Our branching rule is presented in Section V, along with soundness and solution-completeness proofs. Section VI presents experimental results when applying standard CCBS and CCBS using our branching rule to an example problem. Finally, conclusions can be found in VII.
II Background
@@ II-A Problem Formulation
A move action $m=(m_{\varphi},m_{D})$ offers a way for an agent to traverse an edge $(v,v^{\prime})\in\mathcal{E}$. The motion function $m_{\varphi}:[0,m_{D}]\rightarrow\mathcal{M}$ describes a continuous trajectory with duration $m_{D}$ starting at $v$ and ending at $v^{\prime}$; that is, $m_{\varphi}(0)=\mathit{coord}(v)$ and $m_{\varphi}(m_{D})=\mathit{coord}(v^{\prime})$. We denote the source and target vertices as $\mathit{from}(m)=v$ and $\mathit{to}(m)=v^{\prime}$, respectively. Each edge in $\mathcal{E}$ admits a finite number of move actions, providing multiple options for an agent to traverse an edge.
Figure 1 illustrates this concept through edge $(v,v^{\prime})$ and three associated move actions $m_{1}$, $m_{2}$, and $m_{3}$. While all move actions start and end at the same vertices, their trajectories (by $m_{1,\varphi}$, $m_{2,\varphi}$, $m_{3,\varphi}$) and durations ($m_{1,D}$, $m_{2,D}$, $m_{3,D}$) may be distinct.
@@
A timed action $\langle a,t\rangle$ (with $a$ either a move or wait action) specifies that an agent executes $a$ starting at time $t\in\Re$, such that the motion function $a_{\varphi}$ is shifted by $t$. That is, the location in $\mathcal{M}$ of an agent executing $\langle a,t\rangle$ is $a_{\varphi}(t^{\prime}-t)$ during $t^{\prime}\in[t,t+a_{D}]$. In other words, an action specifies how an agent moves and the duration to do so (always starting at time 0), while a timed action additionally specifies when the movement begins.
@@
For a plan $\pi_{i}$ to be valid, it must start at time 0 ($t_{1}^{i}=0$), start and end at $i$’s start and goal vertices ($\mathit{from}(a_{1}^{i})=S(i)$ and $\mathit{to}(a_{n}^{i})=G(i)$), and every successive timed action must start when and where the previous ends ($\forall k=2,..,n$: $t_{k}^{i}=t_{k-1}^{i}+a_{k-1,D}^{i}$ and $\mathit{from}(a_{k}^{i})=\mathit{to}(a_{k-1}^{i})$). Additionally, the final action $a_{n}^{i}$ is an infinite wait action with duration $a_{n,D}^{i}=\infty$ to capture that agents always remain idle indefinitely at their last location. The duration of $\pi_{i}$ is $\pi_{i,D}=\sum_{k=1}^{n-1}a_{k,D}^{i}$, naturally excluding the final infinite wait action.
@@
Figure 2 illustrates an example of a valid plan $\pi_{i}$ starting at $S(i)$ and ending at $G(i)$. The first timed move action is $\langle m_{1}^{i}=\langle m_{1,\varphi}^{i},1.5\rangle,0\rangle$, which specifies that agent $i$ begins at $t=0$ to move according to the motion function $m_{1,\varphi}^{i}$ which has duration $1.5$. This is followed by the timed wait action $\langle w_{2}^{i}=\langle w_{2,\varphi}^{i},0.5\rangle,1.5\rangle$ specifying that $i$ waits at $v_{2}$ from $t=1.5$ for a duration of $0.5$, that is until $t=2$. Thereafter, $i$ executes $\langle m_{3}^{i}=\langle m_{3,\varphi}^{i},1\rangle,2\rangle$ to arrive at $v_{3}$ at $t=3$, and then executes $\langle m_{4}^{i}=\langle m_{4,\varphi}^{i},1.3\rangle,3\rangle$ to arrive at $v_{4}$ at $t=4.3$. Finally, the timed wait action $\langle w_{5}^{i}=\langle w_{5,\varphi}^{i},\infty\rangle,4.3\rangle$ specifies that $i$ waits at $v_{4}$ indefinitely from $t=4.3$. Naturally, the plan $\pi_{i}$ illustrated in the example represents merely one among countless valid plans; alternative wait actions could be employed at any of the vertices, different move actions could be selected for the same edges, or entirely different paths through the graph could be traversed.
@@
Each agent occupies a non-zero volume; a collision occurs when the volumes of two agents overlap. A conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ arises when agents $i$ and $j$ collide at some time while respectively executing $\langle a^{i},t^{i}\rangle$ and $\langle a^{j},t^{j}\rangle$. A conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$ with two move actions is a move-move conflict, and a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with a move and a wait action is a move-wait conflict. By convention, the move action in a move-wait conflict is stored in the first position. Note that a wait-wait conflict can only arise after a move-move or move-wait (assuming agents start in collision-free positions). Therefore, we sufficiently consider only move-move and move-wait conflicts. Two plans $\pi_{i}$ and $\pi_{j}$ conflict if there exists a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$.
@@
A joint plan $\Pi$ is a set containing one valid plan for each agent. A solution is a joint plan where no two plans conflict. An optimal solution $\Pi^{*}$ is a solution that minimizes an objective function $\sigma$ over all solutions. The sum-of-costs (sum over plan durations) and makespan (maximum over plan durations) are the two most common objective functions in the MAPF context. Here, we assume that $\sigma(\Pi)$ is strictly monotonically increasing with respect to the maximum duration over all agent plans, $\max_{\pi\in\Pi}\pi_{D}$, which both the sum-of-costs and makespan satisfy.
II-B Algorithmic Properties
@@
In [20], a sound algorithm returns only conflict-free solutions, while an exact algorithm [^1] In [20], the term optimal is used synonymously with exact. However, an algorithm’s optimality typically refers to its time and space complexity and not to the quality of the solutions that it returns. returns only optimal solutions. These terms are never addressed separately in [20] or here. Therefore, we denote both by sound for simplicity. Also note that completeness guarantees the return of a solution if one exists or the report of no solution otherwise, while the weaker solution completeness only guarantees the former.
II-C Continuous-time Conflict Based Search

### v2

# Optimal Multi-agent Path Finding in Continuous Time

@@ Abstract
Continuous-time Conflict Based-Search (CCBS) has long been viewed as the standard optimal baseline for multi-agent path finding in continuous time (MAPFR), yet recent critiques show that the theoretically described CCBS can fail to terminate on solvable MAPFR problems while the publicly available reference implementation can return sub-optimal solutions. This work presents an analytical framework that yields simple and sufficient conditions under which any CCBS-style algorithm is both sound (returns only optimal solutions) and solution complete (terminates on every solvable MAPFR problem). Investigating the reference CCBS implementation reveals that it violates our sufficient conditions for soundness, with counterexamples demonstrating sub-optimality.
@@
Leveraging the framework, we introduce a branching rule ($\delta$-BR) and prove it restores soundness and termination guarantees. Consequently, the resulting CCBS variant is both sound and solution complete. To our knowledge, this is the first MAPFR solver matching the guarantees of the discrete-time CBS. On a constructed example, CCBS with $\delta$-BR improves sum-of-costs from $10.707$ to $9.000$ ($\approx 16\%$ lower) compared to the reference CCBS implementation. Across benchmarks, the reference CCBS implementation is generally able to find solutions faster than CCBS with $\delta$-BR due to its more aggressive pruning. However, this comes at the cost of occasional sub-optimality and potential non-termination when all solutions are pruned, whereas $\delta$-BR preserves optimality and guarantees termination by design. Because $\delta$-BR largely only affects the branching step, it can be adopted as a drop-in replacement in existing codebases, as we show in our provided implementation. Beyond CCBS, the analytical framework and termination criterion provide a systematic way to evaluate other CCBS-like MAPFR solvers and future extensions, thereby offering tools for rigorous analysis of next-generation MAPFR algorithms.
keywords:
Multi-agent Path Finding , Continuous-time Conflict-based Search , Algorithm soundness and solution completeness , Optimal Path planning
journal: Artificial Intelligence
[…]
[chalmers]organization=Division of Systems and Control, Department of Electrical Engineering, Chalmers University of Technology, city=Göteborg, country=Sweden
1 Introduction
@@
As robotics and automation take an increasingly larger role in society, there is a growing need to move agents from where they are to where they need to be without collision. This is the problem of Multi-agent Path Finding (MAPF), which finds application across diverse domains, from computer games [1] to warehouses [2, 3], factories [4] to airports [5]. Progress toward fast and exact solution methods has been steady for decades, targeting the many variants of MAPF.
@@
In its classical formulation, MAPF is about finding a collision-free path for each agent on a graph from its start vertex to goal vertex. Agents are represented as points and can wait at vertices and traverse edges, with time being discrete, edges taking unit-time to traverse, and an objective function based on arrival times at goal vertices being minimized. This problem is NP-hard [6] and has many variations, comprehensively detailed in, e.g., [7, 8, 9]. Early focus was placed on reducing computation and searching for sub-optimal solutions by, e.g., establishing a priority order among agents or restricting the search to a limited time-window [10, 11]. Shortly thereafter, exact methods such as $M^{*}$ [12], Increasing Cost Tree Search [13] and Conflict Based-Search (CBS) [14] were introduced. These methods cleverly avoid searching the entire solution space — which grows exponentially with the number of agents — by adopting a “lazy” approach of planning agents individually and addressing collisions only as they occur. Of the three, CBS has enjoyed the largest adoption with multiple continuations. To name a few, Meta-Agent CBS [15] and Improved CBS [16] for performance enhancements, Enhanced CBS [17] and CBS-Budget [18] for bounded sub-optimal solutions, Rolling Horizon Collision Resolution [19] for lifelong MAPF, and Multi-Goal CBS [20] for MAPF where agents are assigned multiple goal vertices.
@@
The simplifying assumptions of discrete time and unit-length edges in MAPF bring significant drawbacks to its practical applicability; in real-world settings, environments do not always follow grid-like structures, agent speeds and action times vary, and artificial wait times to ensure that agents move in lockstep inflate the objective value. The continuous-time MAPF variant, MAPFR [21], removes such discrete-time assumptions. In MAPFR, agents can traverse edges at any real-valued time by following one of potentially several motion functions that define a continuous trajectory between two connected vertices. Additionally, agents have volumes such that two agents collide when their volumes overlap. The work in [22] addresses MAPFR for sub-optimal solutions, while [23] finds optimal solutions to a simplified MAPFR problem where agents traverse edges in straight lines at constant speeds. To our knowledge, the only known method to address MAPFR in its entirety for optimal solutions is Continuous-time CBS (CCBS) [21] which claims to be sound (only returns optimal solutions) and solution complete (guaranteed to terminate on any solvable MAPFR problem). Since the introduction of CCBS, numerous continuation works have been published, such as [24] for performance enhancements, $T$-Robust CCBS [25] for solutions that are robust to delays, [26] for Any-Angle MAPF, Continuous-time Prioritize Lifelong Planner [27] for lifelong MAPF, and [28] where CCBS is modified and applied to quadcopter drones.
@@
Despite CCBS’s wide-spread adoption, only recently has a critical assessment in [29] established that CCBS does not fulfill its guarantees. The theoretical description of CCBS does not guarantee termination under the existence of an optimal solution (a condition for solution completeness) and the publicly available implementation does not guarantee that a returned solution is optimal (a condition for soundness). This is shown in [29] theoretically and with counter-examples. Therefore, these results reveal a significant gap between the theory and practice of CCBS and any continuation works that rely on the soundness and solution complete guarantees.
In this work, we develop an analytical framework to aid the analysis of CCBS and find sufficient but not necessary requirements for CCBS to be sound and solution complete. We then show that the publicly available CCBS implementation does not satisfy these requirements, which by itself does not mean that CCBS is not sound and solution complete, but it does complement the results in [29] and our experiments showing unsoundness. Leveraging this framework, we introduce a new branching rule ($\delta$-BR) and prove that CCBS using $\delta$-BR is sound and solution complete. That is, CCBS with $\delta$-BR is guaranteed to return an optimal solution within a finite number of iterations on any solvable MAPFR problem. Our only assumption on the objective function is that it is strictly monotonically increasing with respect to the maximum agent arrival time, thus, we guarantee optimality for sum-of-costs, makespan, and any other objective function with this property. Therefore, to our knowledge, we provide the first provably sound and solution-complete method to the MAPFR problem. With our provided implementation, we show that $\delta$-BR can be adopted as a drop-in replacement in existing CCBS implementations. These results are empirically validated on a constructed counterexample and benchmark problems, showing that the existing CCBS implementation is generally able to find solutions faster than CCBS with $\delta$-BR due to its aggressive search space pruning, however, occasionally returns sub-optimal solutions after removing all optimal solutions. Furthermore, the introduced framework and non-termination criteria provide analytical tools for reasoning about CCBS-like solvers and their extensions, thereby laying a foundation for the next generation of sound and solution complete MAPFR algorithms.
@@
The outline of this article is the following: In Section 2, we formally define the MAPFR problem, soundness, and solution completeness. CCBS is also introduced, along with details surrounding the findings in [29]. The framework for analyzing branching rules and CCBS’s algorithmic properties is presented in Section 3. In Section 4, we apply the framework on CCBS’s implemented branching rule. Our branching rule is presented in Section 5, along with soundness and solution completeness proofs. Section 6 presents empirical results. Finally, conclusions are found in 7.
2 Background
@@ 2.1 Problem Formulation
A move action $m=(m_{\varphi},m_{D})$ offers a way for an agent to traverse an edge $(v,v^{\prime})\in\mathcal{E}$. The motion function $m_{\varphi}:[0,m_{D}]\rightarrow\mathcal{M}$ describes a continuous trajectory with duration $m_{D}$ starting at $v$ and ending at $v^{\prime}$; that is, $m_{\varphi}(0)=\mathit{coord}(v)$ and $m_{\varphi}(m_{D})=\mathit{coord}(v^{\prime})$. We denote the source and target vertices as $\mathit{from}(m)=v$ and $\mathit{to}(m)=v^{\prime}$, respectively. Each edge in $\mathcal{E}$ admits a finite number of move actions, providing multiple options for an agent to traverse an edge. Figure 2 illustrates this concept through edge $(v,v^{\prime})$ and three associated move actions $m_{1}$, $m_{2}$, and $m_{3}$. While all move actions start and end at the same vertices, their trajectories (by $m_{1,\varphi}$, $m_{2,\varphi}$, $m_{3,\varphi}$) and durations ($m_{1,D}$, $m_{2,D}$, $m_{3,D}$) may be distinct.
@@
A timed action $\langle a,t\rangle$ (with $a$ either a move or wait action) specifies that an agent executes $a$ starting at time $t\in\Re$, such that the motion function $a_{\varphi}$ is shifted by $t$. That is, the location in $\mathcal{M}$ of an agent executing $\langle a,t\rangle$ is $a_{\varphi}(t^{\prime}-t)$ during $t^{\prime}\in[t,t+a_{D}]$. In other words, an action specifies how an agent moves and the duration to do so, while a timed action additionally specifies when the movement begins.
@@
For a plan $\pi_{i}$ to be valid, it must start at time 0 ($t_{1}^{i}=0$), start and end at $i$’s start and goal vertices ($\mathit{from}(a_{1}^{i})=S(i)$ and $\mathit{to}(a_{n}^{i})=G(i)$), and every successive timed action must start when and where the previous ends,
$\left(t_{k}^{i}=t_{k-1}^{i}+a_{k-1,D}^{i}\right)\wedge\left(\mathit{to}(a_{k-1}^{i})=\mathit{from}(a_{k}^{i})\right)\quad\forall k=2,..,n.$
Additionally, the final action $a_{n}^{i}$ is an infinite wait action with duration $a_{n,D}^{i}=\infty$ to capture that agents always remain idle indefinitely at their last location. The duration of $\pi_{i}$ is $\pi_{i,D}=\sum_{k=1}^{n-1}a_{k,D}^{i}$, naturally excluding the final infinite wait action.
@@
Figure 2 illustrates an example of a valid plan $\pi_{i}$ starting at $S(i)$ and ending at $G(i)$. The first timed move action is $\langle m_{1}^{i}=\langle m_{1,\varphi}^{i},1.5\rangle,0\rangle$, which specifies that agent $i$ begins at $t=0$ to move according to the motion function $m_{1,\varphi}^{i}$ which has duration $1.5$. This is followed by the timed wait action $\langle w_{2}^{i}=\langle w_{2,\varphi}^{i},0.5\rangle,1.5\rangle$ specifying that $i$ waits at $v(w_{2}^{i})=v_{2}$ from $t=1.5$ for a duration of $0.5$, that is until $t=2$. Thereafter, $i$ executes $\langle m_{3}^{i}=\langle m_{3,\varphi}^{i},1\rangle,2\rangle$ to arrive at $v_{3}$ at $t=3$, and then executes $\langle m_{4}^{i}=\langle m_{4,\varphi}^{i},1.3\rangle,3\rangle$ to arrive at $v_{4}$ at $t=4.3$. Finally, the timed wait action $\langle w_{5}^{i}=\langle w_{5,\varphi}^{i},\infty\rangle,4.3\rangle$ specifies that $i$ waits at $v_{4}$ indefinitely from $t=4.3$. Naturally, the plan $\pi_{i}$ illustrated in the example represents merely one among countless valid plans; alternative wait actions could be employed at any of the vertices, different move actions could be selected for the same edges, or an entirely different path through the graph could be traversed.
@@
Each agent occupies a non-zero volume; a collision occurs when the volumes of two agents overlap. A conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ arises when agents $i$ and $j$ collide at some time while respectively executing $\langle a^{i},t^{i}\rangle$ and $\langle a^{j},t^{j}\rangle$. A conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$ with two move actions is a move-move conflict, and a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with a move and a wait action is a move-wait conflict. By convention, the move action in a move-wait conflict is stored in the first position. Note that a wait-wait conflict can only arise after a move-move or move-wait (assuming agents start in collision-free positions). Therefore, we sufficiently consider only move-move and move-wait conflicts, as a wait-wait conflict cannot exist without one of the other conflict types. Two plans $\pi_{i}$ and $\pi_{j}$ conflict if there exists a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$.
@@
A joint plan $\Pi$ is a set containing one valid plan for each agent, a solution is a joint plan that does not contain any two conflicting plans, and an optimal solution $\Pi^{*}$ is a solution that minimizes an objective function $\sigma$ over all solutions. The sum-of-costs $\sigma(\Pi)=\sum_{\pi\in\Pi}\pi_{D}$ (sum over plan durations) and makespan $\sigma(\Pi)=\max_{\pi\in\Pi}\pi_{D}$ (maximum over plan durations) are the two most common objective functions in the MAPF context. Here, we assume that $\sigma(\Pi)$ is strictly monotonically increasing with respect to the maximum duration over all agent plans in $\Pi$, $\max_{\pi\in\Pi}\pi_{D}$, which both the sum-of-costs and makespan satisfy. Therefore, all proofs and guarantees provided in this work are valid for any objective function satisfying this assumption.
2.2 Algorithmic Properties
@@
In [21], a sound algorithm returns only solutions, while an exact algorithm [^1] In [21], the term optimal is used synonymously with exact. However, an algorithm’s optimality typically refers to its time and space complexity and not to the quality of the solutions that it produces. returns only optimal solutions. These terms are never addressed separately in [21] or here. Therefore, we denote both by sound for simplicity. Also note that completeness guarantees the return of a solution if one exists or the report of no solution otherwise, while the weaker solution completeness only guarantees the former.
2.3 Continuous-time Conflict Based Search